OpenStudy (anonymous):

a body starts falling freely from rest , it covers equal disrances in the 1st 3 seconds and the last second of its motion. For how much time will the body remains in air?

OpenStudy (anonymous):

Here goes the solution. Enjoy it!

OpenStudy (anonymous):

how did you find Vo @CarlosGP ?

OpenStudy (anonymous):

Vo is the speed of the body at t=T-1 With this initial speed we know it has to fall same distance in 1 second than in the first 3 seconds, that is 1/2*g*3^2=9/2g=4.5g, thus in one second Vo*1+1/2*g*1^2=Vo+g/2 and that means Vo+g/2=4.5g, then Vo=4g

OpenStudy (anonymous):

Got it now?